# A Local Projection Stabilised HHO Method for the Oseen Problem

### v1

@@ 3 Wellposedness of Discrete Formulation
Applying the Cauchy-Schwarz inequality on $A_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{z}}_{h})$ along with the equivalence of the norms $\epsilon^{1/2}\left\|\cdot\right\|_{\rm dG}$ and $\left\|\cdot\right\|_{\epsilon,h}$, we get
$A_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{z}}_{h})\leq A_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h})^{1/2}A_{\epsilon,h}(\underline{\boldsymbol{z}}_{h},\underline{\boldsymbol{z}}_{h})^{1/2}\leq CA_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h})^{1/2}\epsilon^{1/2}\left\|\underline{\boldsymbol{z}}_{h}\right\|_{\rm dG}.$ (3.8)
Now, using the definition of the DG norm, relations (3.7) and the fact that $\boldsymbol{z}_{F}=0$, we have
$\epsilon\left\|\underline{\boldsymbol{z}}_{h}\right\|_{\rm dG}^{2}=\sum_{T\in\mathcal{T}_{h}}\Big{(}\epsilon\left\|\nabla\boldsymbol{z}_{T}\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\frac{\epsilon}{h_{F}}\left\|\boldsymbol{z}_{F}-\boldsymbol{z}_{T}\right\|^{2}\Big{)}\leq C\sum_{M\in\mathcal{M}_{h}}\frac{\epsilon}{h_{M}^{2}}\left\|\boldsymbol{z}_{M}\right\|^{2}_{M}.$ (3.9)
@@
Using the definition of $G_{b,T}^{k}$ in the bilinear form $A_{\boldsymbol{b},h}(\cdot,\cdot)$ of (2.3) and applying the integration by parts, we get
@@
Using the definition of the bilinear form $B_{h}$ in (2.10) with an integration by parts and $\boldsymbol{z}_{F}=0$, we have
$B_{h}(\underline{\boldsymbol{z}}_{h},p_{h})=\sum_{T\in\mathcal{T}_{h}}\Big{(}(\boldsymbol{z}_{T},\nabla p_{h})-\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{z}_{F}\cdot\boldsymbol{n},p_{h})_{F}\Big{)}=\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{z}_{T},\nabla_{h}p_{h})_{T}.$ (3.12)
Adding (3.11) and (3.12) we obtain
@@
Using the fact that $\boldsymbol{x}_{M}|_{T}=G_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{T})+\nabla_{h}p_{h}$, the definition of $\boldsymbol{z}_{M}$ and relation (3.6) the first term in the summation of (3) becomes
$\sum_{T\in\mathcal{T}_{h}}(G_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{T})+\nabla_{h}p_{h},\boldsymbol{z}_{T})_{T}$ $=\sum_{M\in\mathcal{M}_{h}}(\boldsymbol{x}_{M},\boldsymbol{z}_{M})_{M}$
@@
Now we estimate the second term in (3). This term can be controlled by applying the triangle inequality and then adding and subtracting the reconstruction $G_{{\boldsymbol{b}_{M}},M}$ with $\gamma_{M}\leq\min\{\tau_{M},\rho_{M}\}$ as follows
@@
$\leq\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|_{M}^{2}+\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|K_{M}(\nabla_{h}p_{h})\right\|_{M}^{2}$
$\leq C\sum_{M\in\mathcal{M}_{h}}\tau_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|_{M}^{2}+C\sum_{M\in\mathcal{M}_{h}}\rho_{M}\left\|K_{M}(\nabla_{h}p_{h})\right\|_{M}^{2}$
$\quad+\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|_{M}^{2}.$
$\leq C\left|\!\left|\!\left|(\underline{\boldsymbol{u}}_{h},p_{h})\right|\!\right|\!\right|^{2}+\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|_{M}^{2}.$ (3.15)
To control the second term of (3) we use the boundedness of the operator $K_{M}$, the inverse inequality, equivalence of the norms $\left\|\cdot\right\|_{\epsilon,h}$ and $\epsilon^{1/2}\left\|\cdot\right\|_{{\rm dG}}$, the estimate (2.6) and $\tau_{M}\left\|b\right\|_{0,\infty,M}\sim h_{M}$, to get
@@
$\leq C\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h})\right\|_{M}^{2}$
$\leq C\sum_{M\in\mathcal{M}_{h}}\tau_{M}h_{M}\left\|\boldsymbol{b}\right\|_{0,\infty}\left\|\boldsymbol{b}\right\|_{1,\infty}\left\|\nabla_{h}\boldsymbol{u}_{T}\right\|_{M}^{2}$
$\leq C\sum_{M\in\mathcal{M}_{h}}\Big{(}\frac{h_{M}^{2}\left\|\boldsymbol{b}\right\|_{1,\infty}}{\epsilon+\sigma h_{M}^{2}}\sum_{T\subset M}(\epsilon+\sigma h_{M}^{2})\left\|\nabla\boldsymbol{u}_{T}\right\|_{T}^{2}\Big{)}$
$\leq C\max_{M\in\mathcal{M}_{h}}\frac{h_{M}^{2}}{\epsilon+\sigma h_{M}^{2}}\Big{(}\left\|\underline{\boldsymbol{u}}_{h}\right\|_{\epsilon,h}^{2}+\left\|\underline{\boldsymbol{u}}_{h}\right\|_{\boldsymbol{b}}^{2}\Big{)}\leq C\omega\left|\!\left|\!\left|(\underline{\boldsymbol{u}}_{h},p_{h})\right|\!\right|\!\right|^{2}.$ (3.16)
Combining (3) and (3) and putting in (3) we obtain
@@
$\leq C(\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\frac{\lvert\boldsymbol{b}\cdot\boldsymbol{n}\rvert}{2}\left\|\boldsymbol{u}_{F}-\boldsymbol{u}_{T}\right\|^{2}_{F})^{1/2}(\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|\boldsymbol{x}_{M}\right\|^{2}_{M})^{1/2}$
@@
Using the definition of the reconstruction $G_{\boldsymbol{b}_{M},M}^{k}$ and (3.7), we have the following estimate
$\left\|\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{z}}_{T})\right\|_{T}^{2}$ $=(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{z}_{T},\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\boldsymbol{z}_{T}))_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}\cdot\boldsymbol{n}(\boldsymbol{z}_{F}-\boldsymbol{z}_{T}),\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\boldsymbol{z}_{T}))_{F}$
@@
The LPS stabilisation term can now be estimated using the last inequality and boundedness of the operator $K_{M}$ with $\gamma_{M}\leq\tau_{M}$ and $\tau_{M}\left\|\boldsymbol{b}\right\|_{0,\infty,M}\sim h_{M}$ as follows
@@
$\leq(\sum_{M\in\mathcal{M}_{h}}\tau_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|^{2}_{M})^{1/2}(\sum_{M\in\mathcal{M}_{h}}\frac{\tau_{M}\left\|\boldsymbol{b}\right\|^{2}_{0,\infty,M}\gamma_{M}^{2}}{h_{M}^{2}}\left\|\boldsymbol{x}_{M}\right\|_{M}^{2})^{1/2}$
@@
$A_{N,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{z}}_{h})$ $=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{u}_{T}-\boldsymbol{u}_{F})\cdot\boldsymbol{n},(\boldsymbol{z}_{T})\cdot\boldsymbol{n})_{F}$
@@
$D_{h}^{k}(\tilde{\underline{\boldsymbol{z}}}_{h})=-p_{h}\quad and\quad\left\|\tilde{\underline{\boldsymbol{z}}}_{h}\right\|_{{\rm dG}}+\left\|\tilde{z}_{h}\right\|\leq C\left\|p_{h}\right\|,$ (3.24)
@@
Using the equivalence of the norms $\left\|\cdot\right\|_{\epsilon,h}$ and $\epsilon^{1/2}\left\|\cdot\right\|_{{\rm dG}}$, the Cauchy–Schwarz inequlity and the bound for $\tilde{\underline{\boldsymbol{z}}}_{h}$ in (3.24), we obtain an estimate for the viscous term as
$A_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\tilde{\underline{\boldsymbol{z}}}_{h})\leq CA_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h})^{1/2}\epsilon^{1/2}\left\|\tilde{\underline{\boldsymbol{z}}}_{h}\right\|_{{\rm dG}}\leq\frac{1}{6}\left\|p_{h}\right\|^{2}+C\left\|\underline{\boldsymbol{u}}_{h}\right\|^{2}_{\epsilon,h}.$ (3.25)
Using the definition of the bilinear form $A_{\boldsymbol{b},T}$ in (2.3), the Cauchy–Schwarz inequality, trace inequality and the fact that $\left\|G_{\boldsymbol{b},T}^{k}(\tilde{\underline{\boldsymbol{z}}}_{T})\right\|_{T}\leq C\left\|\tilde{\underline{\boldsymbol{z}}}_{T}\right\|_{\rm dG}$, we get the estimate for the advection term as
$A_{\boldsymbol{b},T}(\underline{\boldsymbol{u}}_{T},\tilde{\underline{\boldsymbol{z}}}_{T})$ $=-(\boldsymbol{u}_{T},G_{\boldsymbol{b},T}^{k}(\tilde{\underline{\boldsymbol{z}}}_{T}))_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}_{TF}^{\ominus}(\boldsymbol{u}_{F}-\boldsymbol{u}_{T}),(\tilde{\underline{\boldsymbol{z}}}_{F}-\tilde{\underline{\boldsymbol{z}}}_{T}))_{F}+\sigma(\boldsymbol{u}_{T},\tilde{\underline{\boldsymbol{z}}}_{T})_{T}$
$\leq C\left\|\underline{\boldsymbol{u}}_{h}\right\|_{\boldsymbol{b},h}\Big{(}\left\|\tilde{\underline{\boldsymbol{z}}}_{h}\right\|_{\rm dG}+(\sum_{T\in\mathcal{T}_{h}}\left\|\tilde{\underline{\boldsymbol{z}}}_{T}\right\|^{2}_{T})^{1/2}\Big{)}\leq\frac{1}{6}\left\|p_{h}\right\|^{2}+C\left\|\underline{\boldsymbol{u}}_{h}\right\|^{2}_{\boldsymbol{b}}+C\left\|\boldsymbol{u}_{h}\right\|^{2}.$ (3.26)
Using the definition of advection reconstruction for the LPS term and applying the trace and Cauchy–Schwarz inequalities, we obtain $\left\|G_{\boldsymbol{b}_{M},T}^{k}(\tilde{\underline{\boldsymbol{z}}}_{T})\right\|_{T}\leq C\left\|\tilde{\underline{\boldsymbol{z}}}_{T}\right\|_{\rm dG}$. This along with the boundedness of the operator $K_{M}$ and (3.24) yield
$A_{S,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{z}}_{h})\leq C(A_{S,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h}))^{1/2}\left\|p_{h}\right\|.$
@@
$A_{N,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{z}}_{h})\leq C(A_{N,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h}))^{1/2}\left\|p_{h}\right\|.$
@@
$(\epsilon+\sigma)\left\|\boldsymbol{u}_{h}\right\|^{2}$ $\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}\epsilon\left\|\nabla\boldsymbol{u}_{T}\right\|_{T}^{2}+\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\frac{\epsilon}{h_{F}}\int_{F}(\boldsymbol{u}_{F}-\boldsymbol{u}_{T})^{2}\Big{)}+\sigma\left\|\boldsymbol{u}_{h}\right\|^{2}$
@@
Now taking $(\underline{\boldsymbol{v}}_{h},q_{h})=(\underline{\boldsymbol{u}}_{h},p_{h})+\frac{1}{1+8C_{1}+\tilde{C}_{1}}\Big{(}\frac{8}{1+\omega}(\underline{\boldsymbol{z}}_{h},0)+2(\epsilon+\sigma)(\tilde{\underline{\boldsymbol{z}}}_{h},0)\Big{)}$ with $C_{1}$ and $\underline{\boldsymbol{z}}_{h}$ as in (3.5) and $\tilde{C}_{1}$, $\tilde{\underline{\boldsymbol{z}}}_{h}$ as in (3.28) we get
@@
Using the properties of $\underline{\boldsymbol{z}}_{h}$ in (3.6)–(3.7) we have
@@
In a similar manner using (3.24) we have

### v2

@@ 3 Wellposedness of Discrete Formulation
Applying the Cauchy-Schwarz inequality on $A_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{z}}_{h})$ along with the equivalence of the norms $\epsilon^{1/2}\left\|\cdot\right\|_{1,h}$ and $\left\|\cdot\right\|_{\epsilon,h}$, we get
$A_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{z}}_{h})\leq A_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h})^{1/2}A_{\epsilon,h}(\underline{\boldsymbol{z}}_{h},\underline{\boldsymbol{z}}_{h})^{1/2}\leq CA_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h})^{1/2}\epsilon^{1/2}\left\|\underline{\boldsymbol{z}}_{h}\right\|_{1,h}.$ (3.8)
Now, using the definition of the norm, the relations (3.7) and the fact that $\boldsymbol{z}_{F}=0$, we have
$\epsilon\left\|\underline{\boldsymbol{z}}_{h}\right\|_{1,h}^{2}=\sum_{T\in\mathcal{T}_{h}}\Big{(}\epsilon\left\|\nabla\boldsymbol{z}_{T}\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\frac{\epsilon}{h_{F}}\left\|\boldsymbol{z}_{F}-\boldsymbol{z}_{T}\right\|_{F}^{2}\Big{)}\leq C\sum_{M\in\mathcal{M}_{h}}\frac{\epsilon}{h_{M}^{2}}\left\|\boldsymbol{z}_{M}\right\|^{2}_{M}.$ (3.9)
@@
Using the definition of $G_{b,T}^{k}$ in the bilinear form $A_{\boldsymbol{b},h}(\cdot,\cdot)$ of (2.3) and applying the integration by parts, we get (see [1, eqn. (18)])
@@
Using the definition of the bilinear form $B_{h}$ in (2.10) with integration by parts and $\boldsymbol{z}_{F}=0$, we have
$B_{h}(\underline{\boldsymbol{z}}_{h},p_{h})=\sum_{T\in\mathcal{T}_{h}}\Big{(}(\boldsymbol{z}_{T},\nabla p_{h})-\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{z}_{F}\cdot\boldsymbol{n}_{TF},p_{h})_{F}\Big{)}=\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{z}_{T},\nabla_{h}p_{h})_{T}.$ (3.12)
Adding (3.11) and (3.12), we obtain
@@
Using the fact that $\boldsymbol{x}_{M}|_{T}=G_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{T})+\nabla p_{T}$, the definition of $\boldsymbol{z}_{M}$ and relation (3.6) the first term in the summation of (3) becomes
$\sum_{T\in\mathcal{T}_{h}}(G_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{T})+\nabla p_{T},\boldsymbol{z}_{T})_{T}$ $=\sum_{M\in\mathcal{M}_{h}}(\boldsymbol{x}_{M},\boldsymbol{z}_{M})_{M}$
@@
Now we estimate the second term in (3). This term can be controlled by applying the triangle inequality and then adding and subtracting the reconstruction $G_{{\boldsymbol{b}_{M}},M}$ with $\gamma_{M}\leq\min\{\tau_{M},\rho_{M}\}$ as follows:
@@
$\leq 2\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|_{M}^{2}+2\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|K_{M}(\nabla_{h}p_{h})\right\|_{M}^{2}$
$\leq 2\sum_{M\in\mathcal{M}_{h}}\tau_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|_{M}^{2}+2\sum_{M\in\mathcal{M}_{h}}\rho_{M}\left\|K_{M}(\nabla_{h}p_{h})\right\|_{M}^{2}$
$\quad+\sum_{M\in\mathcal{M}_{h}}2\gamma_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|_{M}^{2}.$
$\leq 2\left|\!\left|\!\left|(\underline{\boldsymbol{u}}_{h},p_{h})\right|\!\right|\!\right|^{2}+\sum_{M\in\mathcal{M}_{h}}2\gamma_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|_{M}^{2}.$ (3.15)
To control the second term of (3), we use the boundedness of the operator $K_{M}$, the inverse inequality, the equivalence of the norms $\left\|\cdot\right\|_{\epsilon,h}$ and $\epsilon^{1/2}\left\|\cdot\right\|_{1,h}$, the estimate (2.6) and $\tau_{M}\left\|b\right\|_{0,\infty,M}\lesssim h_{M}$, to get
@@
$\leq\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h})\right\|_{M}^{2}$
$\leq C\sum_{M\in\mathcal{M}_{h}}\tau_{M}h_{M}\left\|\boldsymbol{b}\right\|_{0,\infty,M}\left\|\boldsymbol{b}\right\|_{1,\infty,M}\left\|\nabla_{h}\boldsymbol{u}_{h}\right\|_{M}^{2}$
$\leq C\sum_{M\in\mathcal{M}_{h}}\Big{(}\frac{h_{M}^{2}\left\|\boldsymbol{b}\right\|_{1,\infty,M}}{\epsilon+\sigma h_{M}^{2}}\sum_{T\subset M}(\epsilon+\sigma h_{M}^{2})\left\|\nabla\boldsymbol{u}_{T}\right\|_{T}^{2}\Big{)}$
$\leq C\max_{M\in\mathcal{M}_{h}}\frac{h_{M}^{2}\left\|\boldsymbol{b}\right\|_{1,\infty,M}}{\epsilon+\sigma h_{M}^{2}}\Big{(}\left\|\underline{\boldsymbol{u}}_{h}\right\|_{\epsilon,h}^{2}+\left\|\underline{\boldsymbol{u}}_{h}\right\|_{\boldsymbol{b}}^{2}\Big{)}\leq C\omega\left|\!\left|\!\left|(\underline{\boldsymbol{u}}_{h},p_{h})\right|\!\right|\!\right|^{2}.$ (3.16)
Combining (3) and (3) and putting in (3), we obtain
@@
$\leq C(\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\left\|\boldsymbol{b}_{TF}(\boldsymbol{u}_{F}-\boldsymbol{u}_{T})\right\|^{2}_{F})^{1/2}(\sum_{M\in\mathcal{M}_{h}}\gamma_{M}\left\|\boldsymbol{x}_{M}\right\|^{2}_{M})^{1/2}$
@@
Using the definition of reconstruction $G_{\boldsymbol{b}_{M},M}^{k}$ and (3.7), we have the following estimate
$\left\|\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{z}}_{T})\right\|_{T}^{2}$ $=(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{z}_{T},\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\boldsymbol{z}_{T}))_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}\cdot\boldsymbol{n}_{TF}(\boldsymbol{z}_{F}-\boldsymbol{z}_{T}),\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\boldsymbol{z}_{T}))_{F}$
@@
The LPS stabilisation term can now be estimated using the last inequality and boundedness of the operator $K_{M}$ with $\gamma_{M}\leq\tau_{M}$ and $\tau_{M}\left\|\boldsymbol{b}\right\|_{0,\infty,M}\lesssim h_{M}$ as follows
@@
$\leq\Big{(}\sum_{M\in\mathcal{M}_{h}}\tau_{M}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{u}}_{h}))\right\|^{2}_{M}\Big{)}^{1/2}\Big{(}\sum_{M\in\mathcal{M}_{h}}\frac{\tau_{M}\left\|\boldsymbol{b}\right\|^{2}_{0,\infty,M}\gamma_{M}^{2}}{h_{M}^{2}}\left\|\boldsymbol{x}_{M}\right\|_{M}^{2}\Big{)}^{1/2}$
@@
$A_{N,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{z}}_{h})$ $=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{u}_{T}-\boldsymbol{u}_{F})\cdot\boldsymbol{n}_{TF},(\boldsymbol{z}_{T})\cdot\boldsymbol{n}_{TF})_{F}$
@@
$D_{h}^{k}(\tilde{\underline{\boldsymbol{z}}}_{h})=-p_{h}\quad and\quad\left\|\tilde{\underline{\boldsymbol{z}}}_{h}\right\|_{1,h}\leq C\left\|p_{h}\right\|,$ (3.24)
@@
Using the equivalence of the norms $\left\|\cdot\right\|_{\epsilon,h}$ and $\epsilon^{1/2}\left\|\cdot\right\|_{1,h}$, the Cauchy–Schwarz inequlity, and the bound for $\tilde{\underline{\boldsymbol{z}}}_{h}$ in (3.24), we obtain an estimate for the viscous term as
$A_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\tilde{\underline{\boldsymbol{z}}}_{h})\leq CA_{\epsilon,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h})^{1/2}\epsilon^{1/2}\left\|\tilde{\underline{\boldsymbol{z}}}_{h}\right\|_{1,h}\leq\frac{1}{6}\left\|p_{h}\right\|^{2}+C\left\|\underline{\boldsymbol{u}}_{h}\right\|^{2}_{\epsilon,h}.$ (3.25)
Using the definition of the bilinear form $A_{\boldsymbol{b},T}$ in (2.3), the Cauchy–Schwarz inequality, the discrete Poincaré inequality (1.3) and the fact that $\left\|G_{\boldsymbol{b},T}^{k}(\tilde{\underline{\boldsymbol{z}}}_{T})\right\|_{T}\leq C\left\|\tilde{\underline{\boldsymbol{z}}}_{T}\right\|_{1,h}$, we get the estimate for the advection term as
$A_{\boldsymbol{b},T}(\underline{\boldsymbol{u}}_{T},\tilde{\underline{\boldsymbol{z}}}_{T})$ $=-(\boldsymbol{u}_{T},G_{\boldsymbol{b},T}^{k}(\tilde{\underline{\boldsymbol{z}}}_{T}))_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}_{TF}^{\ominus}(\boldsymbol{u}_{F}-\boldsymbol{u}_{T}),(\tilde{\boldsymbol{z}}_{F}-\tilde{\boldsymbol{z}}_{T}))_{F}+\sigma(\boldsymbol{u}_{T},\tilde{\underline{\boldsymbol{z}}}_{T})_{T}$
$\leq C(\left\|\underline{\boldsymbol{u}}_{h}\right\|_{\boldsymbol{b}}+\|u_{h}\|)(\left\|\tilde{\underline{\boldsymbol{z}}}_{h}\right\|_{1,h}+\|\tilde{\boldsymbol{z}}_{h}\|)\leq\frac{1}{6}\left\|p_{h}\right\|^{2}+C(\left\|\underline{\boldsymbol{u}}_{h}\right\|^{2}_{\boldsymbol{b}}+\left\|\boldsymbol{u}_{h}\right\|^{2}).$ (3.26)
Using the definition of advection reconstruction for the LPS term and applying the trace and Cauchy–Schwarz inequalities, we obtain $\left\|G_{\boldsymbol{b}_{M},T}^{k}(\tilde{\underline{\boldsymbol{z}}}_{T})\right\|_{T}\leq C\left\|\tilde{\underline{\boldsymbol{z}}}_{T}\right\|_{1,h}$. This, along with the boundedness of the operator $K_{M}$ and (3.24) yields
$A_{S,h}(\underline{\boldsymbol{u}}_{h},\tilde{\underline{\boldsymbol{z}}}_{h})\leq C(A_{S,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h}))^{1/2}\left\|p_{h}\right\|.$
@@
$A_{N,h}(\underline{\boldsymbol{u}}_{h},\tilde{\underline{\boldsymbol{z}}}_{h})\leq C(A_{N,h}(\underline{\boldsymbol{u}}_{h},\underline{\boldsymbol{u}}_{h}))^{1/2}\left\|p_{h}\right\|.$
@@
$(\epsilon+\sigma)\left\|\boldsymbol{u}_{h}\right\|^{2}$ $\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}\epsilon\left\|\nabla\boldsymbol{u}_{T}\right\|_{T}^{2}+\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\frac{\epsilon}{h_{F}}\int_{F}(\boldsymbol{u}_{F}-\boldsymbol{u}_{T})\cdot(\boldsymbol{u}_{F}-\boldsymbol{u}_{T})\Big{)}+\sigma\left\|\boldsymbol{u}_{h}\right\|^{2}$
@@
Now taking $(\underline{\boldsymbol{v}}_{h},q_{h})=(\underline{\boldsymbol{u}}_{h},p_{h})+\frac{1}{1+8C_{1}+\tilde{C}_{1}}\Big{(}\frac{8}{1+\omega}(\underline{\boldsymbol{z}}_{h},0)+2(\epsilon+\sigma)(\tilde{\underline{\boldsymbol{z}}}_{h},0)\Big{)}$ with $C_{1}$ and $\underline{\boldsymbol{z}}_{h}$ as in (3.5) and $\tilde{C}_{1}$, $\tilde{\underline{\boldsymbol{z}}}_{h}$ as in (3.28), we get
@@
Using the properties of $\underline{\boldsymbol{z}}_{h}$ in (3.6)–(3.7), we have
@@
In a similar manner using (3.24), we have